OpenStudy (anonymous):

The value of ΔG°' for the conversion of glucose-6-phosphate to fructose-6-phosphate (F6P) is 1.67 kJ/mol. If the concentration of glucose-6-phosphate at equilibrium is 2.35 mM, what is the concentration of fructose-6-phosphate? Assume a temperature of 25.0 °C. The constant R = 8.3145 J/(mol·K)

OpenStudy (aaronq):

First find eq. constant, K \(\sf \large \Delta G=-RTlnK\rightarrow K=e^{-\frac{\Delta G}{RT}}\) The equation is a rearragement: \(G6P \rightleftharpoons F6P\) \(\sf \large K=\dfrac{[F6P]}{[G6P]}\rightarrow [G6P]=K*[F6P]\) \(\sf \large [G6P]= e^{-\frac{\Delta G}{RT}}*[F6P]\)

OpenStudy (aaronq):

oh is solved for 6GP not F6P as i should have. just rearrange the equation